Find Rate for Investment. Assume you invest $500 in a savings certificate offering an interest rate that is compounded annually to reach a maturity value of $1,000 at the end of 6 years. What is the interest rate offered on this investment?

A) 8.08%

B) 12.24%

C) 4.59%

D) 13.24%

E) 16.30%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1000=500*(1+r/100)^6

(1000/500)^(1/6)=(1+r/100)

(1+r/100)=1.1224

r=1.1224-1

=12.24%(Approx).